OpenStudy (anonymous):

let f(x)=1-2x^2 and g(x) = x+1 compute the following (f o f)(x)

OpenStudy (anonymous):

do you mean (f o g)(x) ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do I time the f(x) twice

OpenStudy (anonymous):

\[f\circ f(x)=f(f(x))=f(1-2x^2)=1-2(1-2x^2)^2\]

OpenStudy (anonymous):

satellite he meant f o g not f o f

OpenStudy (anonymous):

no I mean f 0 f that is correct

OpenStudy (anonymous):

oh then satellite has it

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

so if i have (g o g) do I go like this

OpenStudy (anonymous):

but use the formula for g?

OpenStudy (anonymous):

to find gog, substitute x with the equation of g in the formula of g. gog=g(g(x))=g(x+1)=(x+1)+1=x+2